# ShadowLLM: Predictor-based Contextual Sparsity for Large Language Models

### v1

@@ Abstract
The high power consumption and latency-sensitive deployments of large language models (LLMs) have motivated techniques like quantization and sparsity. Contextual sparsity, where the sparsity pattern is input-dependent, is crucial in LLMs because the permanent removal of attention heads or neurons from LLMs can significantly degrade accuracy. Prior work has attempted to model contextual sparsity using neural networks trained to predict activation magnitudes, which can be used to dynamically prune structures with low predicted activation magnitude. In this paper, we look beyond magnitude-based pruning criteria to assess attention head and neuron importance in LLMs. We developed a novel predictor called ShadowLLM, which can shadow the LLM behavior and enforce better sparsity patterns, resulting in over 15% improvement in end-to-end accuracy without increasing latency compared to previous methods. ShadowLLM achieves up to a 20% speed-up over the state-of-the-art DejaVu framework. These enhancements are validated on models with up to 30 billion parameters. Our code is available at ShadowLLM.
@@
Yash Akhauri1,2, Ahmed F AbouElhamayed1, Jordan Dotzel1,2, Zhiru Zhang1, Alexander M Rush1, Safeen Huda2, and Mohamed S Abdelfattah1 1Cornell University 2Google {ya255, afa55, jad443}@cornell.edu {zhiruz, arush, mohamed}@cornell.edu,  safeen@google.com
Figure 1: ShadowLLM uses more accurate pruning criteria and a simpler sparsity predictor compared to DejaVu. Its pruning criteria results in a stronger accuracy-sparsity trade-off across seven downstream evaluation tasks, and its unified predictor improves the execution latency compared to the layerwise predictor of DejaVu.
@@ 1 Introduction
A key optimization in LLM deployment is sparsification, where weights or activations are pruned to reduce the computation and memory requirements at run time. Sparsification can either be static, which permanently removes an attention head or neuron, or contextual, which prunes based on the current input and context. While some works investigate task-specific static pruning methods Bansal et al. (2022); Michel et al. (2019), their impact on in-context learning abilities reduce accuracy compared to contextual sparsity.
@@ 2.1 Pruning Criteria
Research in discovering good criteria for pruning neurons has focused on using the activations, weights, and gradients of neural networks to assess the relative importance of neurons. Several pruning criteria have been designed to utilize light-weight computations, such as a single forward-backward pass through the network. For instance, some works use parameter magnitudes as a proxy for parameter saliency Frankle and Carbin (2018); Han et al. (2015), whereas others use gradient-based information LeCun et al. (1989); Hassibi and Stork (1992); Molchanov et al. (2016); Bansal et al. (2022). Further, research in Neural Architecture Search (NAS) adapts these pruning criteria to assess and compare entire architectures. Such initialization-based measures like NASWOT Mellor et al. (2021) aim to study other properties of the architecture, and can be used to study neuron importance as well.
@@ 2.2 LLM Inference Optimization
Within these works, DejaVu leverages dynamic sparsity by building predictors to estimate sparsity patterns. In this paper, we investigate how the predictor can be improved, both in terms of performance and preserving model quality.
@@ 3 Pruning Criteria
Figure 6: Global pruning outperforms local (per-layer) pruning strategies using ShadowLLM trained on the plainact criteria. Global pruning accommodates the varying importance of different layers, allowing for unbalanced pruning across layers.
@@
Among these methods, we find that a gradient-based sensitivity method we call plainact out-performs activation-based magnitude methods adapted in prior dynamic pruning research Liu et al. (2023). The L2Norm activation-magnitude based criterion assesses the importance of neurons by simply taking the L2 Norm of the head and neuron activation as $||A_{l,k}||_{2}$. The plainact criterion measures the expected sensitivity of the model on the loss if a head or neuron is removed. For the head and neuron, this can be described simply as $||A_{l,k}\cdot\frac{\partial\mathcal{L}}{\partial A_{l,k}}||_{1}$ and $||\theta_{l,k}\cdot\frac{\partial\mathcal{L}}{\partial\theta_{l,k}}||_{1}$ respectively. We perform in-depth ablations across several pruning criteria in Section 6, and find that plainact is a good pruning criteria.
@@ 4 Predictors For Neuron Ranking
When deploying a large language model, for a given input, we will not have access to the activations or the gradients. Thus, calculating the L2Norm or plainact criterion is not possible. However, it is possible to create a dataset of inputs and their corresponding L2Norm or plainact for each head and neuron. Such a dataset can be used to train a predictor, which can take the input and predict the sparsity pattern of the model at deployment. Sparsity prediction can reduce the end-to-end latency of transformer inference by predicting which operations can be skipped.
@@
Table 1: For a transformer with E embedding dimension, N layers, H heads, F FFN neurons per layer, ShadowLLM uses (N-1)Ep1 fewer FLOPs, where p1 is the predictor hidden dimension. The table also shows the percentage improvement in predictor FLOPs for ShadowLLM vs DejaVu for different models.
@@
DejaVu employs a two-layer MLP, taking the activation from the final token at every alternating layer and predicting the sparsity of the next layer. A significant portion of the complexity of the DejaVu system arises from its asynchronous look-ahead predictor, which is implemented to hide the predictor latency. While DejaVu aggressively optimizes an asynchronous look-ahead predictor, these predictors can be expensive. The predictor itself only takes 2% of the total FLOPs for OPT-1.3B, but having a per-layer predictor adds significant overhead due to additional kernel launches and memory bandwidth constraints, leading to an end-to-end latency increase of 25% over static sparsity (same sparsity but fixed, without a predictor).
@@
DejaVu’s per-layer approach to pruning, where a fixed sparsity is enforced per layer, can be sub-optimal as true contextual sparsity should be independent of layers. To study our proposed predictor designs in a contextual-sparsification setting, we evaluate the Spearman-$\rho$ rank coefficient between the relative importance order of neurons and heads given by the predictor, and the relative importance order given by the pruning criterion. Additionally, this is done on a global head-ranking task. From Figure 6, we see that DejaVu-style layer-wise predictors are not trained for global pruning. We find that Full Seq. ShadowLLM performs similarly to ShadowLLM, but with a significant increase in overall FLOPs.
@@
Figure 7: Comparison of the DejaVu-style predictor trained on a magnitude based metric (L2Norm) with ShadowLLM on the best pruning criteria (plainact) on WikiText2. In both local and global settings, ShadowLLM performs well due to better pruning criteria.
@@
To analyze the ability of ShadowLLM predictors to assess neuron importance in a global and local (per-layer) setting, we train ShadowLLM predictors for plainact across all seven down-stream tasks in the 5-shot setting, with a per-layer (local) output normalization scheme. We then evaluate the WikiText2 perplexity of the OPT-1.3B model as we increase sparsity for both the global and local (per-layer) pruning strategies. For local pruning, every layer achieves the target sparsity, and relative importance are only compared intra-layer. In Figure 6, we find that ShadowLLM is able to preserve perplexity in both global and local cases. However, we find that global pruning generally performs better than per-layer pruning. This can be attributed to the fact that some layer heads are more important than others, and forcing over-parameterization of a layer (due to uniform pruning across every layer) may cause a more important head to be pruned out from an under-parameterized layer.
@@ 5 Evaluation
We find that the activation-gradient based pruning criteria that we use in ShadowLLM are effective for downstream evaluation tasks as well as perplexity. Further, we demonstrated in Section 4 that ShadowLLM can predict the sparsity pattern for the entire LLM given just the input to the first layer. We find that the accuracy and predictability trade-off is excellent for the plainact criterion, whereas other criteria were harder to learn due to outliers and high variance. In this section, we evaluate the effectiveness of combining ShadowLLM predictor design with the plainact criterion, compared to our implementation of DejaVu-style [^1] To enable comparisons across pruning criteria, we have implemented our own DejaVu-style predictor in ShadowLLM predictors, trained on a magnitude based pruning criteria.
@@ 5.1 Experimental Setup
Our ablation studies to identify good pruning criteria, as well as test the efficacy of predictors is conducted on OPT-1.3B. Further, local and global pruning strategies are tested on OPT-13B and OPT-30B. Our downstream evaluation across seven tasks is reported on OPT-13B.
@@ 5.2 Model Quality
In Figure 7, we train the DejaVu-style and ShadowLLM predictors on their respective pruning criterion (L2Norm and plainact respectively) on 3000 input-output examples across 7 downstream tasks in the five-shot setting. The perplexity is evaluated in a local and global pruning setting on WikiText2.
Global pruning enables better model quality - sparsity trade-off. Thus, in Figure 8 we evaluate OPT-13B by training the ShadowLLM and DejaVu-style predictors in the same setting, and doing downstream evaluation in the zero-shot setting across seven tasks. We show that there is a consistent accuracy improvement across tasks, attributed to better pruning criteria.
@@ 5.3 Performance
We implement the ShadowLLM predictor along with the prior enhancements introduced by DejaVu and conduct our performance experiments on up to 4 A100 GPUs. In Figure 10, we measure the end-to-end time taken for a generation length of 128 tokens at 50% sparsity and observe an average 16.07% improvement over DejaVu. Figure 12 shows a consistent improvement in generation time as output tokens increase. Further, Figure 12 shows that ShadowLLM is on average 21.25% faster than DejaVu. Finally, we profile model sizes from 1.3B to 66B, observing up to a 13.95% improvement in time per-inference.
@@
Figure 16: The aggregate importance score per neuron is used to conduct static pruning of the LLM when testing perplexity on WikiText2. fisher and plainact preserve model quality better than other criteria.
@@ 6 Analysis
In Figure 14, we calculate the fisher criteria for every neuron and head on 3500 input-output examples from the downstream tasks for the 0-shot, 3-shot, and 5-shot settings. We average the criteria for each neuron and head across these examples and evaluate WikiText2 perplexity as model sparsity is increased. The results indicate that providing more in-context examples when registering the criteria improves model quality during pruning.
@@
In Figure 16, we use the task pruning criterion averaged over their respective examples for each head and neuron to do a static sparsification of the OPT-1.3B model and test it in the 0-shot setting. We report the mean accuracy across the downstream tasks for each pruning criteria. We find that jacov is an stable criteria to preserve model performance in the static case. However, jacov does not have a context-dependent equivalent, as it relies on the covariance of Jacobian matrices across examples. We evaluate these proxies in Figure 16, and find that fisher and plainact preserve model quality well, with jacov performing worse. jacov might have higher task-dependence for static pruning, and does not translate to better general model quality.
@@
While we can shadow activation magnitudes with a predictor, we need to balance finding the best pruning criteria for assessing neuron importance with ensuring the criteria is easy to learn. To identify such a criteria, we measure each criteria for each head and neuron on $\approx$3000 input-output examples across the 7 downstream tasks in a 5-shot setting. We train our predictor to use the output of the first attention layer’s last sequence index to predict per-head and neuron importance. Figure 14 reports the average Spearman-$\rho$ rank correlation on $\approx$ 400 input-output examples. From Figure 16, we see that fisher delivers the best perplexity for up to 50% sparsity, but delivers a Spearman-$\rho$ of under 0.7. Similarly, grasp is difficult to predict due to its high range and outliers. In contrast, we find that the plainact criterion is easy to predict and performs well in a contextual setting.
@@ 7 Conclusion
In this paper, we present an improved pruning criterion for contextual sparsity, that leverages both activations and gradients, leading to significant improvements in model quality. We demonstrate that these criteria can be effectively modeled by a single predictor at the first layer of the LLM, eliminating the need for calculation at deployment. Our findings, validated on models with up to 30 billion parameters, show that relatively small predictors can effectively model contextual sparsity in LLMs. Using gradient-informed pruning criteria, we can accurately identify important neurons and heads across various model sizes. Our approach of combining an enhanced pruning criterion with an early predictor, results in over 15% improvement in accuracy without increasing latency, and a 20% improvement in performance across different model sizes.
@@ A.1 Predictor Design
Input Dimension OPT Model Embedding
@@ A.2 Additional Pruning Criteria
Figure 18: Each pruning criterion is measured and averaged per neuron and head over 3500 training examples in a 5-shot setting from all downstream tasks. This aggregate importance score per neuron and head is used to conduct static pruning of the LLM when testing perplexity on WikiText2. fisher and plainact preserve model quality better than other criteria.

### v2

@@ Abstract
The high power consumption and latency-sensitive deployments of large language models (LLMs) have motivated efficiency techniques like quantization and sparsity. Contextual sparsity, where the sparsity pattern is input-dependent, is crucial in LLMs because the permanent removal of attention heads or neurons from LLMs can significantly degrade accuracy. Prior work has attempted to model contextual sparsity using neural networks trained to predict activation magnitudes, which can be used to dynamically prune structures with low predicted activation magnitude. In this paper, we look beyond magnitude-based pruning criteria to assess attention head and neuron importance in LLMs. We develop a novel predictor called ShadowLLM, which can shadow the LLM behavior and enforce better sparsity patterns, resulting in over 15% improvement in end-to-end accuracy compared to prior methods. In addition, ShadowLLM achieves up to a 20% speed-up over the state-of-the-art DejaVu framework. These enhancements are validated on Llama-2 and OPT models with up to 30 billion parameters. Our code is available at ShadowLLM.
@@
Yash Akhauri1,2, Ahmed F. AbouElhamayed1, Jordan Dotzel1,2, Zhiru Zhang1, Alexander M. Rush1, Safeen Huda2, and Mohamed S. Abdelfattah1 1Cornell University 2Google {ya255, afa55, jad443}@cornell.edu {zhiruz, arush, mohamed}@cornell.edu,  safeen@google.com
Figure 1: ShadowLLM uses more accurate pruning criteria and a simpler sparsity predictor compared to DejaVu. Its pruning criteria results in a stronger accuracy-sparsity trade-off (geomean) across seven downstream evaluation tasks, and its unified predictor improves the execution latency compared to the layerwise predictor of DejaVu.
@@ 1 Introduction
A key optimization in LLM deployment is sparsification, where weights or activations are pruned to reduce the computation and memory requirements at run time. Sparsification can either be static, which permanently removes an attention head or neuron, or contextual, which prunes based on the current input and context. While some works investigate task-specific static pruning methods Bansal et al. (2022); Michel et al. (2019), they typically have a large impact on in-context learning, reducing downstream task accuracy compared to contextual sparsity.
@@ 2.1 Pruning Criteria
Research in the area of designing criteria for pruning neurons has focused on using the activations, weights, and gradients of neural networks to assess the relative importance of neurons. Several pruning criteria have been designed to utilize light-weight computations, such as a single forward-backward pass through the network. For instance, some works use parameter magnitudes as a proxy for parameter saliency Frankle and Carbin (2018); Han et al. (2015), whereas others use gradient-based information LeCun et al. (1989); Hassibi and Stork (1992); Molchanov et al. (2016); Bansal et al. (2022). Further, research in Neural Architecture Search (NAS) adapts these pruning criteria to assess and compare entire architectures. Such initialization-based measures like NASWOT Mellor et al. (2021) aim to study other properties of the architecture, and can be used to study neuron importance as well.
@@ 2.2 LLM Inference Optimization
Within these works, DejaVu Liu et al. (2023) leverages dynamic sparsity by building predictors to estimate sparsity patterns. In this paper, we investigate how the predictor can be improved, both in terms of performance and model quality.
@@ 3 Pruning Criteria
Figure 6: Global pruning outperforms local (per-layer) pruning strategies using ShadowLLM trained on the plainact criteria (OPT-1.3B). Global pruning accommodates the varying importance of different layers, allowing for unbalanced pruning across layers.
@@
Among these methods, we find that a gradient-based sensitivity method we call plainact out-performs activation-based magnitude methods adapted in prior dynamic pruning research Liu et al. (2023). The L2Norm activation-magnitude based criterion assesses the importance of neurons by simply taking the L2 Norm of the head and neuron activation as $||A_{l,k}||_{2}$. The plainact criterion measures the expected sensitivity of the model on the loss if a head or neuron is removed. For the head and neuron, this can be described simply as $||A_{l,k}\cdot\frac{\partial\mathcal{L}}{\partial A_{l,k}}||_{1}$ and $||\theta_{l,k}\cdot\frac{\partial\mathcal{L}}{\partial\theta_{l,k}}||_{1}$ respectively. We perform in-depth ablations across several pruning criteria in Section 6, and find that plainact empirically performs well as a pruning criteria.
@@ 4 Predictors For Neuron Ranking
When deploying a large language model, for a given input, we will not have access to the activations or the gradients. Thus, calculating the L2Norm or plainact criterion is not possible. However, it is possible to create a calibration dataset of inputs and their corresponding L2Norm or plainact for each head and neuron. Such a dataset can be used to train a predictor, which can take the input and predict the sparsity pattern of the model at deployment. Sparsity prediction can reduce the end-to-end latency of transformer inference by predicting which operations can be skipped.
@@
Table 1: For a transformer with E embedding dimension, N layers, H heads, F FFN neurons per layer, ShadowLLM uses (N-1)Ep1 fewer FLOPs, where p1 is the predictor hidden dimension. The table also shows the percentage improvement in predictor FLOPs for ShadowLLM vs. DejaVu for different models.
@@
DejaVu employs a two-layer MLP, taking the activation from the final token at every alternating layer and predicting the sparsity of the next layer. A significant portion of the complexity of the DejaVu system arises from its asynchronous look-ahead predictor which can be expensive in wall clock time despite aggressive optimizations within DejaVu. The predictor itself only takes 2% of the total FLOPs for OPT-1.3B, but having a per-layer predictor adds significant overhead due to additional GPU kernel launches and memory bandwidth constraints, leading to an end-to-end latency increase of 25% over static sparsity (same sparsity but fixed, without a predictor).
@@
DejaVu’s per-layer approach to pruning, where a fixed sparsity is enforced per layer, can be sub-optimal as true contextual sparsity should be independent of layers. To study our proposed predictor designs in a contextual-sparsification setting, we evaluate the Spearman-$\rho$ (rank correlation coefficient) between the relative importance order of neurons and heads given by the predictor, and the relative importance order given by the pruning criterion. Additionally, this is done on a global head-ranking task. From Figure 6, we see that DejaVu-style layer-wise predictors are not trained for global pruning. We find that Full Seq. ShadowLLM performs similarly to ShadowLLM, but with a significant increase in overall FLOPs.
@@
Figure 7: Comparison of the DejaVu-style predictor trained on a magnitude-based metric (L2Norm) with ShadowLLM on the best pruning criteria (plainact) on WikiText2. In both local and global settings, ShadowLLM performs well due to better pruning criteria.
@@
To analyze the ability of ShadowLLM predictors to assess neuron importance in a global and local (per-layer) setting, we train ShadowLLM predictors for plainact across all seven down-stream tasks in the 5-shot setting, with a per-layer (local) output normalization scheme. We then evaluate the WikiText2 perplexity  [^1] For effective context-sparsity evaluation, perplexity calculations are performed on a per-document basis, differing from standard concatenation methods; see Section B for details. of the OPT-1.3B model as we increase sparsity for both the global and local (per-layer) pruning strategies. For local pruning, every layer achieves the target sparsity, and relative importance are only compared intra-layer. In Figure 6, we find that ShadowLLM is able to preserve perplexity in both global and local cases. However, we find that global pruning generally performs better than per-layer pruning. This can be attributed to the fact that some layer heads are more important than others, and forcing equal pruning ratios for all layers may cause over-parameterization in some layers and more important head to be pruned out from an under-parameterized layer.
@@ 5 Evaluation
We find that the activation-gradient based pruning criteria that we use in ShadowLLM are effective for downstream evaluation tasks as well as perplexity. Further, we demonstrated in Section 4 that ShadowLLM can predict the sparsity pattern for the entire LLM given just the input to the first layer. We find that the accuracy and predictability trade-off is excellent for the plainact criterion, whereas other criteria were harder to learn due to outliers and high variance. In this section, we evaluate the effectiveness of combining ShadowLLM predictor design with the plainact criterion, compared to our implementation of DejaVu-style [^2] To enable comparisons across pruning criteria, we have implemented our own DejaVu-style predictor in ShadowLLM. predictors, trained on a magnitude based pruning criteria.
@@
Figure 9: Gradient-informed criteria (Plainact) improves global pruning on Llama-2-7b, resulting in an end-to-end perplexity improvement on WikiText2.
@@ 5.1 Experimental Setup
Our ablation studies to identify good pruning criteria, as well as test the efficacy of predictors is conducted on OPT-1.3B. Further, local and global pruning strategies are tested on OPT-13B and OPT-30B, and global pruning on Llama-2-7b Touvron et al. (2023). Our downstream evaluation across seven tasks is reported on OPT-13B.
@@ 5.2 Model Quality
In Figure 7, we train the DejaVu-style and ShadowLLM predictors on their respective pruning criterion (L2Norm and plainact respectively) on 2720 input-output examples across 7 downstream tasks in the five-shot setting. The perplexity is evaluated in a local and global pruning setting on WikiText2.
Global pruning enables better model quality - sparsity trade-off. Figure 10 compares the perplexity-sparsity trade-off on Llama-2-7b model, with Plainact significantly improving perplexity. Further, in Figure 8 we evaluate OPT-13B by training the ShadowLLM and DejaVu-style predictors in the same setting, and doing downstream evaluation in the zero-shot setting across seven tasks. We show that there is a consistent accuracy improvement across tasks, attributed to better pruning criteria.
@@ 5.3 Performance
We implement the ShadowLLM predictor along with the prior enhancements introduced by DejaVu and conduct our performance experiments on up to 4 A100 GPUs. In Figure 10, we measure the end-to-end time taken for a generation length of 128 tokens at 50% sparsity and observe an average 16.2% improvement over DejaVu. Figure 12 shows a consistent improvement in generation time as output tokens increase. Further, Figure 12 shows that ShadowLLM is on average 21.25% faster than DejaVu in the decode phase specifically. Finally, we profile model sizes from 1.3B to 66B, observing up to a 21.3% improvement in time per-inference.
@@
Figure 16: The aggregate importance score per neuron is used to conduct static pruning of the LLM. fisher and plainact preserve model quality better than other criteria. Dashed black line is dense baseline.
@@ 6 Analysis
In Figure 14, we calculate the fisher criteria for every neuron and head on 2720 input-output examples from the downstream tasks for the 0-shot, 3-shot, and 5-shot settings. We average the criteria for each neuron and head across these examples and evaluate WikiText2 perplexity as model sparsity is increased. The results indicate that providing more in-context examples when registering the criteria improves model quality during pruning.
@@
In Figure 16, we use the task pruning criterion averaged over their respective examples for each head and neuron to do a static sparsification of the OPT-1.3B model and test it in the 0-shot setting. We report the mean accuracy across the downstream tasks for each pruning criteria. We find that jacov is a stable criteria to preserve model performance in the static case. However, jacov does not have a context-dependent equivalent, as it relies on the covariance of Jacobian matrices across examples. We evaluate these proxies in Figure 16, and find that fisher and plainact preserve model quality well, with jacov performing worse. jacov might have higher task-dependence for static pruning, and does not translate to better general model quality.
@@
While we can shadow activation magnitudes with a predictor, we need to balance finding the best pruning criteria for assessing neuron importance, ensuring the criteria is easy to learn. To identify such a criteria, we measure each criteria for each head and neuron on 2720 input-output examples across the 7 downstream tasks in a 5-shot setting. We train our predictor to use the output of the first attention layer’s last sequence index to predict per-head and neuron importance. Figure 14 reports the average Spearman-$\rho$ rank correlation on 680 input-output examples. From Figure 16, we see that fisher delivers the best perplexity for up to 50% sparsity, but delivers a Spearman-$\rho$ of under 0.7. Similarly, grasp is difficult to predict due to its high range and outliers. In contrast, we find that the plainact criterion is easy to predict and performs well in a contextual setting.
@@ 7 Conclusion
In this paper, we work towards significantly simplifying the predictor design, and study several pruning criteria. However, our study is limited to smaller models, up to 30B parameters on only OPT style models. Further, criteria like nwot are designed for the ReLU activation function, which may not be directly applicable to attention maps. We profile these for completeness regardless, however, more research in pruning criteria is needed. Finally, we train predictors on less than 10000 input-output examples, more examples may enable better sparsity pattern modeling.
Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. 2339084, in addition to funding by Intel Corporation. We would like to thank Nilesh Jain, Juan Pablo Munoz, Sameh Gobriel, and Vui Seng Chua for helpful discussions and feedback.
@@ A.1 Predictor Design
Input Dimension Model Embedding
@@ A.2 Additional Pruning Criteria
Figure 19: Each pruning criterion is measured and averaged per neuron and head over 2720 training examples in a 5-shot setting from all downstream tasks. This aggregate importance score per neuron and head is used to conduct static pruning of the LLM when testing perplexity on WikiText2. fisher and plainact preserve model quality better than other criteria.
@@
Model Ours Reported
Lin et al. (2024)
Llama2-7B 8.82 5.47
OPT-1.3B 16.4 14.6
OPT-13B 11.5 10.1
OPT-30B 10.7 9.56
Table 4: Comparison of our perplexity and reported perplexity for various models. Following Bansal et al. (2022), we calculate per-document perplexity, which increases model perplexity.
Appendix B On Perplexity Calculation
[…]